# Sparsity-Promoting Reachability Analysis and Optimization of Constrained Zonotopes

### v1

@@ Abstract
The constrained zonotope is a polytopic set representation widely used for set-based analysis and control of dynamic systems. This paper considers the problem of tailoring a quadratic program (QP) optimization algorithm to the particular structure of constrained zonotopes and vice-versa. An alternating direction method of multipliers (ADMM) algorithm is presented that makes efficient use of the constrained zonotope structure. To increase the efficiency of the ADMM iterations, reachability calculations are presented that increase the sparsity of the matrices used to define a constrained zonotope. Numerical results show that the ADMM algorithm solves optimal control problems built using these reachability calculations faster than state-of-the-art QP solvers using conventional problem formulations, especially for large problems. Constrained zonotope reachability and optimization calculations are combined within a set-valued state estimation and moving horizon estimation algorithm, and a projection-based infeasibility detection method is presented for efficient safety verification of system trajectories.
@@ I Introduction
Novel set-theoretic methods continue to be developed as important tools for analyzing and controlling dynamic systems. Sets can be used to rigorously bound the reachable states of a system, for example, or to define constraints to which a controller must adhere. Zonotopes and generalizations of zonotopes have become particularly popular set representations, as is evidenced by their appearance of several open-source toolboxes [1, 2, 3, 4, 5, 6, 7]. Key to the appeal of these set representations is that they have closed-form expressions for many important set operations and, in the context of reachability analysis, have favorable complexity growth when compared to traditional set representations such as halfspace representation (H-rep) polytopes, vertex representation (V-rep) polytopes, and ellipsoids [8].
@@
While it is typically very efficient to perform operations on constrained zonotopes, analyzing these sets often requires the use of numerical optimization algorithms. For example, assessing whether a constrained zonotope contains a point requires solving a linear program (LP) feasibility problem [9]. Similarly, assessing whether a constrained zonotope is empty also requires solving a feasibility problem. By contrast, point containment for H-rep polytopes only requires matrix-vector multiplication and inequality evaluation. Despite this fundamental reliance on numerical optimization, the problem of tailoring optimization algorithms to constrained zonotopes has received little attention.
@@ I-A Gaps in the Literature
In the specific context of motion planning, the particular properties and structure of hybrid zonotopes are exploited within a mixed-integer optimization routine in [24, 25]. While some of the optimization techniques presented in these papers apply to constrained zonotopes, the presented methods are restricted to multi-stage problem structures and problems with diagonal cost matrices. Structure exploitation for more general optimization problems formulated using constrained zonotopes has not been considered.
@@ I-B Contributions
This paper considers the problems of exploiting the structure of constrained zonotopes within an optimization algorithm and developing reachability calculations to produce set representations that can be more efficiently analyzed by that optimization algorithm. The key contributions of this paper are as follows: (1) Novel reachability calculations for affine dynamics using constrained zonotopes decrease the number of non-zero elements in the matrices that define the set when compared to existing methods. (2) An efficient ADMM formulation solves convex quadratic programs (QPs) for which the constraint set is represented as a constrained zonotope. (3) A method is proposed for efficiently verifying that a constrained zonotope is empty using the ADMM iterations. (4) An algorithm based on reachability analysis of constrained zonotopes combines moving horizon estimation (MHE) and SVSE.
@@ I-C Outline
The remainder of this paper is organized as follows. Preliminary information is given in Sec. II. In Sec. III, reachability calculations for linear systems using constrained zonotopes are presented that result in matrices with high sparsity. Sec. IV proposes an ADMM algorithm that is tailored to constrained zonotopes. In Sec. V, the sparse reachability calculations and ADMM algorithm are applied to MPC, MHE, and safety verification problems. Numerical results for these problems are given in Sec. VI. Using the proposed approach based on sparse constrained zonotope reachability analysis and ADMM, the MPC problems are solved faster than when using state-of-the-art QP solvers and a commonly-used sparse MPC problem formulation, especially for large problem sizes. The MHE example elucidates an equivalence between reachability analysis and MHE as a set-valued state estimate is extracted directly from the MHE problem formulation. The infeasibility certificate is used for efficient safety verification of system trajectories in a third numerical example. Sec. VII concludes the paper.
@@ II-A Notation
Unless otherwise stated, scalars are denoted by lowercase letters, vectors by boldface lowercase letters, matrices by uppercase letters, and sets by calligraphic letters. $\mathbf{0}=\begin{bmatrix}0&\cdots&0\end{bmatrix}^{T}$ and $\mathbf{1}=\begin{bmatrix}1&\cdots&1\end{bmatrix}^{T}$ denote vectors consisting entirely of zeroes and ones, respectively, of appropriate dimensions. $I$ denotes the identity matrix. The operator $\operatorname*{\mathrm{nnz}}{(\cdot)}$ gives the number of non-zero elements of a sparse matrix. Empty brackets [][][ ] denote an appropriately-dimensioned matrix where one or both dimensions are zero. For a set $\mathcal{C}$ and a function $f(\mathbf{x})$, $\min_{\mathbf{x}\in\mathcal{C}}f(\mathbf{x})$ and $\max_{\mathbf{x}\in\mathcal{C}}f(\mathbf{x})$ denote the minimum and maximum values of $f(\mathbf{x})$ such that $\mathbf{x}\in\mathcal{C}$. The indicator function for a set $\mathcal{C}$ is defined as
$I_{\mathcal{C}}(\mathbf{x})=\begin{cases}{0,\;\mathbf{x}\in\mathcal{C}}\;,\\ {\infty,\;\text{otherwise}}\;.\end{cases}$
A closed interval $[x]$ is defined by its lower bound $\underline{x}$ and upper bound $\overline{x}$, as $[x]=[\underline{x},\overline{x}]$. A box $[\mathbf{x}]$ is a vector of closed intervals such that $[\mathbf{x}]=\begin{bmatrix}[x_{1}]&[x_{2}]&\cdots\end{bmatrix}^{T}$.
@@ II-B Zonotopes and Constrained Zonotopes
II-C Alternating Direction Method of Multiers
@@
The algorithm is defined by the iterations
$\mathbf{x}_{k+1}=\operatorname*{\mathrm{argmin}}_{\mathbf{x}}(f(\mathbf{x})+\frac{\rho}{2}||A\mathbf{x}+B\mathbf{z}_{k}-\mathbf{c}+\mathbf{u}_{k}||_{2}^{2})\;,$ (5a)
$\mathbf{z}_{k+1}=\operatorname*{\mathrm{argmin}}_{\mathbf{z}}(g(\mathbf{z})+\frac{\rho}{2}||A\mathbf{x}_{k+1}+B\mathbf{z}-\mathbf{c}+\mathbf{u}_{k}||_{2}^{2})\;,$ (5b)
@@
$\mathbf{r}_{p}=A\mathbf{x}_{k}+B\mathbf{z}_{k}-\mathbf{c}\;,\mathbf{r}_{d}=\rho A^{T}B(\mathbf{z}_{k+1}-\mathbf{z}_{k})\;.$ (6)
Common absolute convergence criteria are $||\mathbf{r}_{p}||_{2}<\sqrt{p}\epsilon_{p}$ and $||\mathbf{r}_{d}||_{2}<\sqrt{n}\epsilon_{d}$ with $\epsilon_{p}>0$ and $\epsilon_{d}>0$ the primal and dual convergence tolerances, respectively [19].
@@ II-D Interval Arithmetic
$\alpha[x]=\begin{cases}[\alpha\underline{x},\alpha\overline{x}],\;\text{if }\alpha\geq 0\;,\\ [\alpha\overline{x},\alpha\underline{x}],\;\text{if }\alpha<0\;,\end{cases}$ (7a)
$[x]+[y]=[\underline{x}+\underline{y},\overline{x}+\overline{y}]\;,$ (7b)
$[x]-[y]=[\underline{x}-\overline{y},\overline{x}-\underline{y}]\;.$ (7c)
@@
[…]
@@ III Reachability with Constrained Zonotopes via Sparsity-Promoting Operations
[…]
@@ III-A1 Standard Method
$A_{x,k+1}=\begin{bmatrix}A_{x,k}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ -AG_{x,k}&-BG_{u}&G_{s}\end{bmatrix}\;.$ (11b)
@@ III-A2 Graph of Function Method
In [29], a reachability analysis technique using graphs of functions was proposed. While that work targeted nonlinear and hybrid systems, graphs of functions can be used to perform reachability calculations for affine systems as well. In the current context, the graph of a function $\Psi$ is given as
@@
$\Psi=\left\langle\begin{bmatrix}G_{s}&0&0\\ 0&G_{u}&0\\ AG_{s}&BG_{u}&0\end{bmatrix},\cdot\begin{bmatrix}A_{s}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ -AG_{s}&-BG_{u}&G_{s}\end{bmatrix},\cdot\right\rangle\;.$ (14)
Defining the rows of the $G$ matrix in (14) to be $G_{\psi,x,k}$, $G_{\psi,u,k}$, and $G_{\psi,x,k+1}$ and the columns of the $A$ matrix to be $A_{\psi,x,k}$, $A_{\psi,u,k}$, and $A_{\psi,x,k+1}$, the generator and constraint matrices for the iterates (13) are given explicitly as
@@
$A_{x,k+1}=\begin{bmatrix}A_{\psi,x,k}&A_{\psi,u,k}&A_{\psi,x,k+1}&0&0\\ 0&0&0&A_{x,k}&0\\ 0&0&0&0&A_{u}\\ -G_{\psi,x,k}&0&0&G_{x,k}&0\\ 0&-G_{\psi,u,k}&0&0&G_{u,k}\end{bmatrix}\;.$ (15b)
@@ III-A3 Sparsity-Promoting Method
Eq. (16) gives the set of reachable states for the linear system (9) at time step $k+1$ given the set $\mathcal{X}_{k}$ of possible states at time step $k$.
@@
$A_{x,k+1}=\begin{bmatrix}A_{x,k}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ -AG_{x,k}&-BG_{u}&G_{s}\end{bmatrix}\;.$ (17b)
@@ III-B Numerical Example: Reachable Sets of a Second-Order System
[…]
Figure 1: Reachable sets $\mathcal{X}_{k}$ for the second order system (18). The gray set is the $N=15$-step reachable set $\mathcal{X}_{N}$.
@@
TABLE I: Number of non-zero elements in the generator matrix $G$ and constraint matrix $A$ for $\mathcal{X}_{N}$.
@@ IV-A ADMM Formulation
$\min_{\bm{\xi}}\frac{1}{2}\bm{\xi}^{T}G^{T}PG\bm{\xi}+[G^{T}(P\mathbf{c}+\mathbf{q})]^{T}\bm{\xi}\;,$ (20a)
[…]
@@
it is clear that (20) is a QP with constraints $\bm{\xi}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}=\left\{\bm{\xi}\in^{n_{G}}\middle|A\bm{\xi}=\mathbf{b}\right\}$ is an affine set and =$\mathcal{B}=\left\{\bm{\xi}\in^{n_{G}}\middle|-\mathbf{1}\leq\bm{\xi}\leq\mathbf{1}\right\}$ is a box. Using interval notation, we will alternately write $\mathcal{B}$ as $[\bm{\beta}]$.
@@
If Assumption 1 does not hold, then redundant constraints can be removed as in [12, Alg. 1].
@@
Note that $M$ will not have fewer non-zero elements using a constrained zonotope representation $\mathcal{Z}$ vs. H-rep in general. In practice, $M$ has fewer non-zero elements when 1) sparsity-promoting operations are used to compute $\mathcal{Z}$ (Sec. III), and 2) the number of equality constraints in $\mathcal{Z}$ is minimized (e.g., by using operations on zonotopes to build $\mathcal{Z}$ when possible).

### v2

@@ Abstract
The constrained zonotope is a polytopic set representation widely used for set-based analysis and control of dynamic systems. This paper develops methods to formulate and solve optimization problems for dynamic systems in real time using constrained zonotope reachability analysis. An alternating direction method of multipliers (ADMM) algorithm is presented that makes efficient use of the constrained zonotope structure. To increase the efficiency of the ADMM iterations, reachability calculations are presented that increase the sparsity of the matrices used to define a constrained zonotope when compared to typical methods. The developed methods are used to formulate and solve predictive control, state estimation, and safety verification problems. Numerical results show that optimization times using the proposed approach are competitive with state-of-the-art QP solvers and conventional problem formulations. A combined set-valued state estimation and moving horizon estimation algorithm is presented and experimentally demonstrated in the context of robot localization.
@@ I Introduction
Novel set-theoretic methods continue to be developed as important tools for analyzing and controlling dynamic systems. For example, sets can be used to rigorously bound the reachable states of a system or to define constraints to which a controller must adhere. Zonotopes and generalizations of zonotopes have become particularly popular set representations, as is evidenced by their appearance in several open-source toolboxes [1, 2, 3, 4, 5, 6, 7, 8]. Key to the appeal of these set representations is that they have closed-form expressions for many important set operations and, in the context of reachability analysis, have favorable complexity growth when compared to traditional set representations such as halfspace representation (H-rep) polytopes, vertex representation (V-rep) polytopes, and ellipsoids [9].
These attributes make zonotopes and generalizations of zonotopes attractive candidates for use in real-time reachability analysis. In [10], for example, zonotope-based reachability analysis was used to verify the safety of system trajectories in autonomous driving experiments. In addition to safety verification, reachability analysis can be used to formulate optimal control and planning problems, as in [11, 12, 13, 14].
@@
While it is typically very efficient to perform operations on constrained zonotopes, analyzing these sets often requires the use of numerical optimization algorithms. For example, assessing whether a constrained zonotope contains a point requires solving a linear program (LP) feasibility problem [15]. Similarly, assessing whether a constrained zonotope is empty also requires solving a feasibility problem. By contrast, point containment for H-rep polytopes only requires matrix-vector multiplication and inequality evaluation. Despite this fundamental reliance on numerical optimization, the problem of constructing these sets to be efficiently analyzed by specific optimization algorithms has received little attention.
@@ I-A Gaps in the Literature
In the specific contexts of motion planning and moving horizon estimation, the particular properties and structure of hybrid zonotopes are exploited within a mixed-integer optimization routine in [30, 31, 32]. While some of the optimization techniques presented in these papers apply to constrained zonotopes, the methods are restricted to multi-stage problem structures and problems with diagonal cost matrices. Efficient solution of more general optimization problems formulated using constrained zonotopes has not been considered.
@@ I-B Contributions
This paper considers the real-time formulation and solution of constrained optimization problems for dynamic systems. Our approach is illustrated in Fig. 1, where the feasible space of a constrained optimal control or estimation problem for a discrete-time dynamic system is constructed using constrained zonotope reachability analysis, and then ADMM is used to find the optimal trajectory within the reachable sets. Expressing the feasible space of an optimization problem through set-theoretic operations provides a unique framework from which to consider its structure and computational complexity. This leads to new insights in constructing a problem representation that is amenable to efficient optimization, and in selecting an optimization algorithm to most efficiently exploit the problem representation. We pursue computational efficiencies in both the reachability and optimization steps to facilitate their real-time execution.
The key contributions of this paper are as follows: (1) Novel reachability calculations for affine dynamics using constrained zonotopes decrease the number of non-zero elements in the matrices that define the set when compared to existing methods. (2) An ADMM formulation that makes efficient use of the constrained zonotope structure solves convex quadratic programs (QPs) for which the constraint set is represented as a constrained zonotope. (3) A method is proposed for efficiently verifying that a constrained zonotope is empty using the ADMM iterations. (4) An algorithm based on reachability analysis of constrained zonotopes combines moving horizon estimation (MHE) and SVSE. The effectiveness of this algorithm is demonstrated experimentally.
@@
112233445566011223344Reachable setsOptimal trajectory
Figure 1: Online reachability analysis with constrained zonotopes is used to construct the feasible space of an optimal control or estimation problem. ADMM is then used to find a trajectory within the reachable sets that minimizes a specified quadratic objective function.
@@ I-C Outline
The remainder of this paper is organized as follows. Preliminary information is given in Sec. II. In Sec. III, reachability calculations for linear systems using constrained zonotopes are presented that result in matrices with high sparsity. Sec. IV proposes an ADMM algorithm that is tailored to constrained zonotopes. In Sec. V, the sparse reachability calculations and ADMM algorithm are applied to MPC, MHE, and safety verification problems. Numerical results for these problems are given in Sec. VI. Using the proposed approach based on sparse constrained zonotope reachability analysis and ADMM, the MPC problems are solved faster than when using state-of-the-art QP solvers and a commonly used sparse MPC problem formulation, especially for large problem sizes. The SVSE/MHE example illustrates an equivalence between reachability analysis and MHE, as a set-valued state estimate is extracted directly from the MHE problem formulation. The infeasibility certificate is used for efficient safety verification of system trajectories in a third numerical example. Sec. VII experimentally demonstrates the real-time capabilities of the proposed methods, with the combined SVSE/MHE algorithm applied for online robot localization. Sec. VIII concludes the paper.
@@ II-A Notation
Unless otherwise stated, scalars are denoted by lowercase letters, vectors by boldface lowercase letters, matrices by uppercase letters, and sets by calligraphic letters. $\mathbf{0}=\begin{bmatrix}0&\cdots&0\end{bmatrix}^{T}$ and $\mathbf{1}=\begin{bmatrix}1&\cdots&1\end{bmatrix}^{T}$ denote vectors consisting entirely of zeroes and ones, respectively, of appropriate dimensions. $I$ denotes the identity matrix. The operator $\operatorname*{\mathrm{nnz}}{(\cdot)}$ gives the number of non-zero elements of a sparse matrix. For a set $\mathcal{C}$ and a function $f(\mathbf{x})$, $\min_{\mathbf{x}\in\mathcal{C}}f(\mathbf{x})$ and $\max_{\mathbf{x}\in\mathcal{C}}f(\mathbf{x})$ denote the minimum and maximum values of $f(\mathbf{x})$ such that $\mathbf{x}\in\mathcal{C}$. The indicator function for a set $\mathcal{C}$ is defined as
$I_{\mathcal{C}}(\mathbf{x})=\begin{cases}{0,\;\text{ if }\mathbf{x}\in\mathcal{C}}\;,\\ {\infty,\;\text{otherwise}}\;.\end{cases}$
A closed interval $[x]$ is defined by its lower bound $\underline{x}$ and upper bound $\overline{x}$, as $x]=[\underline{x},\overline{x}]$. A box $\mathbf{x}]$ is a vector of closed intervals such that $\mathbf{x}]=\begin{bmatrix}[x_{1}]&[x_{2}]&\cdots\end{bmatrix}^{T}$.
@@ II-B Zonotopes and Constrained Zonotopes
II-C Alternating Direction Method of Multipliers
@@
$\text{s.t.}\;A\mathbf{x}+B\mathbf{z}=\mathbf{c}\;.$ (4b)
The algorithm is defined by the iterations
@@
[…]
@@
$\mathbf{r}_{p}=A\mathbf{x}_{k}+B\mathbf{z}_{k}-\mathbf{c}\;,\quad\mathbf{r}_{d}=\rho A^{T}B(\mathbf{z}_{k+1}-\mathbf{z}_{k})\;.$ (6)
Common absolute convergence criteria are $\mathbf{r}_{p}||_{2}<\sqrt{p}\epsilon_{p}$ and $\mathbf{r}_{d}||_{2}<\sqrt{n}\epsilon_{d}$ with $\epsilon_{p}>0$ and $\epsilon_{d}>0$ the primal and dual convergence tolerances, respectively [25].
@@ II-D Interval Arithmetic
$[x]+[y]=[\underline{x}+\underline{y},\overline{x}+\overline{y}]\;,$ (7a)
$[x]-[y]=[\underline{x}-\overline{y},\overline{x}-\underline{y}]\;,$ (7b)
$\alpha[x]=\begin{cases}[\alpha\underline{x},\alpha\overline{x}],\;\text{if }\alpha\geq 0\;,\\ [\alpha\overline{x},\alpha\underline{x}],\;\text{if }\alpha<0\;,\end{cases}$ (7c)
$[x][y]=[\min{(\underline{x}\underline{y},\underline{x}\overline{y},\overline{x}\underline{y},\overline{x}\overline{y})},\max{(\underline{x}\underline{y},\underline{x}\overline{y},\overline{x}\underline{y},\overline{x}\overline{y})}]\;.$ (7d)
@@
A scalar $x\in[x]$ if and only if $\underline{x}\leq x\leq\overline{x}$. Similarly, a vector $\mathbf{x}\in[\mathbf{x}]$ if and only if $x_{i}\in[x_{i}]\;\forall i$.
@@ III Reachability with Constrained Zonotopes via Sparsity-Promoting Operations
$\text{s.t.}\;\forall k:\mathbf{x}_{k},\mathbf{x}_{k+1}\in\mathcal{S}\;,\;\mathbf{u}_{k}\in\mathcal{U}\;,$ (9b)
@@ III-A1 Standard Method
$A_{x,k+1}=\begin{bmatrix}A_{x,k}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ AG_{x,k}&BG_{u}&-G_{s}\end{bmatrix}\;$. (11b)
@@ III-A2 Graph of Function Method
In [36], a reachability analysis technique using graphs of functions was proposed. The graph of the function $\psi(\cdot)$ is defined as
$\Psi=\left\{\begin{bmatrix}\mathbf{p}&\mathbf{q}\end{bmatrix}^{T}\middle|\mathbf{p}\in\mathcal{D},\mathbf{q}\in\psi(\mathbf{q})\subseteq\mathcal{Q}\right\}\;,$ (12)
where $\mathcal{D}$ is the domain set of the inputs $\mathbf{p}$ and $\mathcal{Q}$ is the domain set of the outputs $\mathbf{q}$. While that work targeted nonlinear and hybrid systems, graphs of functions can be used to perform reachability calculations for affine systems as well. For the system in (9), the graph of the function $\psi(\mathbf{x},\mathbf{u})=A\mathbf{x}+B\mathbf{u}$ with $\mathcal{D}=\mathcal{S}\times\mathcal{U}$ and $\mathcal{Q}=\mathcal{S}$ is given as
@@
$\Psi=\left\langle\begin{bmatrix}G_{s}&0&0\\ 0&G_{u}&0\\ AG_{s}&BG_{u}&0\end{bmatrix},\cdot,\begin{bmatrix}A_{s}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ AG_{s}&BG_{u}&-G_{s}\end{bmatrix}$$\cdot\right\rangle\;$. (15)
Defining the rows of the generator matrix in (15) to be $G_{\psi,x,k}$, $G_{\psi,u,k}$, and $G_{\psi,x,k+1}$ and the columns of the constraint matrix to be $A_{\psi,x,k}$, $A_{\psi,u,k}$, and $A_{\psi,x,k+1}$, the generator and constraint matrices for the iterates (14) are given explicitly as
@@
$A_{x,k+1}=\begin{bmatrix}A_{\psi,x,k}&A_{\psi,u,k}&A_{\psi,x,k+1}&0&0\\ 0&0&0&A_{x,k}&0\\ 0&0&0&0&A_{u}\\ G_{\psi,x,k}&0&0&-G_{x,k}&0\\ 0&G_{\psi,u,k}&0&0&-G_{u,k}\end{bmatrix}\;$. (16b)
@@ III-A3 Sparsity-Promoting Method
Eq. (17) gives the one-step reachable set of system (9).
@@
$A_{x,k+1}=\begin{bmatrix}A_{x,k}&0&0\\ 0&A_{u}&0\\ 0&0&A_{s}\\ AG_{x,k}&BG_{u}&-G_{s}\end{bmatrix}\;$. (18b)
@@ III-B Numerical Example: Reachable Sets of a Second-Order System
0.25-0.250.000.000.250.250.500.500.750.751.001.00xx0.5-0.50.00.00.50.51.01.0vv
Figure 2: Reachable sets $\mathcal{X}_{k}$ for the second-order linear system (19). The gray set is the $N=15$-step reachable set $\mathcal{X}_{N}$.
@@
TABLE I: Number of non-zero elements in the generator matrix $G$ and constraint matrix AA for $\mathcal{X}_{N}$.
@@ IV-A ADMM Formulation
$\min_{\bm{\xi}}\frac{1}{2}\bm{\xi}^{T}G^{T}PG\bm{\xi}+\left(G^{T}(P\mathbf{c}+\mathbf{q})\right)^{T}\bm{\xi}\;,$ (21a)
$\text{s.t.}\;A\bm{\xi}=\mathbf{b},\;\bm{\xi}\in[-1,1]^{n_{G}}\;,$ (21b)
@@
it is clear that (21) is a QP with constraints $\bm{\xi}\in\mathcal{A}\cap\mathcal{B}$ where $\mathcal{A}=\left\{\bm{\xi}\in^{n_{G}}\middle|A\bm{\xi}=\mathbf{b}\right\}$ is an affine set and $\mathcal{B}=\left\{\bm{\xi}\in^{n_{G}}\middle|-\mathbf{1}\leq\bm{\xi}\leq\mathbf{1}\right\}$ is a box. Using interval notation, we will alternately write $\mathcal{B}$ as $\bm{\beta}]$.
@@
If Assumption 1 does not hold, then redundant constraints can be removed as in [18, Alg. 1].
@@
Note that $M$ will not have fewer non-zero elements using a constrained zonotope representation $\mathcal{Z}$ vs. H-rep in general. In practice, $M$ has fewer non-zero elements when (1) sparsity-promoting operations are used to compute $\mathcal{Z}$ (Sec. III), and (2) the number of equality constraints in $\mathcal{Z}$ is minimized (e.g., by using operations on zonotopes to build $\mathcal{Z}$ when possible).